OpenStudy (anonymous):

hey guys i need to find all values of x for which the tangent of f(x)=x-sinx is horizontal

OpenStudy (anonymous):

would the derivative be 1-2cos(2x)?

OpenStudy (freckles):

how did you see which to double angle?

OpenStudy (freckles):

switch*

OpenStudy (anonymous):

oh i made a typo in the question it should be f(x)=x-sin2x

OpenStudy (freckles):

So yeah you did good on the derivative part \[1-2\cos(2x)=0 \\ -2\cos(2x)=-1 \\ \cos(2x)=\frac{1}{2}\]

OpenStudy (freckles):

now find when cos(u)=1/2 on the unit circle

OpenStudy (anonymous):

pi/3 and 2pi/3

OpenStudy (anonymous):

should i multiply them by 2? because of the 2x?

OpenStudy (freckles):

pi/3 and 5pi/3

OpenStudy (anonymous):

true 5pi/3

OpenStudy (freckles):

\[\text{ so we had } \cos(u)=\frac{1}{2} \text{ and found } \\ u=\frac{\pi}{3}+2\pi n \text{ or } u=\frac{5 \pi}{3}+2n \pi \text{ but } u =2x\]

OpenStudy (freckles):

\[2x=\frac{\pi}{3}+2\pi n \text{ or } 2x=\frac{5 \pi}{3}+2 n \pi \]

OpenStudy (freckles):

to find x you divide both sides by 2

OpenStudy (anonymous):

so it would be pi/6+2pi(n) or 5pi/6+2pi(n)